# A Multi-Component, Multi-Physics Computational Model for Solving Coupled Cardiac Electromechanics and Vascular Haemodynamics

### v1

@@ 3 Results and Discussion
In this section, we present and analyse the results of the three simulation tests in Sections 3.13.3. Lastly, we assess the computational efficiency of the coupled model. All our simulations are performed on the ARCHER2 UK national supercomputer (https://www.archer2.ac.uk).
3.1 Test 0: Left Ventricle
We perform the standalone simulation of Alya for two heartbeats. We present the results in the second heartbeat since the electromechanical model usually requires some beats to reach a periodic solution. The simulation results are shown in Figure 6, presenting the time series of cardiac pressure and volume as well as the displacement of heart muscles at the end of the ejection phase. Additionally, the endocardial pressure immediately before the ejection phase in the second heartbeat suggests a reference pressure $P_{\text{ref}}$ of 68 mmHg for the fluid domain in the subsequent tests.
(a) Left ventricular pressure and volume
(b) Displacement of heart muscles
Figure 6: (a) Left ventricular pressure and volume and (b) displacement of heart muscles at the end of the ejection phase ($t=0.97$ s) in the standalone simulation of Alya. The left ventricular pressure immediately before the ejection phase suggests a reference pressure $P_{\text{ref}}$ of 68 mmHg for the fluid domain in the subsequent tests.
3.2 Test 1: Left Ventricle with Cylinder
Here we present the results of the coupled simulations using the left ventricle and cylinder geometries. First of all, we assess the reliability of the coupling scheme. To test the implementation of the coupling, we perform the coupled simulations using different time step sizes and coupling frequencies in HemeLB while keeping them fixed in Alya. These simulations are completed at the specified simulation times, showing that our implementation of the coupling algorithm is reliable.
Figure 7: Pressure and flow rate values used by Alya and HemeLB at the coupling interface. The results show that the values are consistent in the two codes during coupling.
Second, we check the consistency of the exchanged data. In the following, we focus on the simulation with $k=4$. The pressure and flow rate values output by Alya and HemeLB at the coupling interface are shown in Figure 7. During the coupling phase, the outputs agree very well for pressure and flow rate. This suggests the data exchanged are consistent in the two solvers and that our implementation of the coupling algorithm is correct.
@@ 3.2 Test 1: Left Ventricle with Cylinder
Figure 8: (a) Pressure and (b) flow rate on the inlet, middle, and outlet planes of the cylinder domain. The variables are influenced by wave reflection and the viscous sponge layer around the outlet.
Third, we evaluate the stability of the coupled simulation. The pressure and flow rate at the inlet, middle, and outlet of the fluid domain are plotted in Figure 8. As shown in the figure, these flow variables remain bounded throughout the simulation. The same is observed for the solutions of Alya. We found that the WK2 model and the viscous sponge layer at the fluid outlet effectively dampen numerical oscillations in the flow, thereby stabilising the simulation.
3.3 Test 2: Left Ventricle with Thoracic Aorta
Here we present the results of the simulations using the left ventricle and thoracic aorta geometries. Initially, the coupled simulations become unstable at the tested coupling frequencies ($k=8,10,12,14,16,20$). However, applying the low-pass filter with an appropriate smoothing factor ($\alpha$), all these simulations become stable. This demonstrates the effectiveness of the low-pass filter in stabilising the coupled model for this specific system. In the following, we look into the simulation using $k=16$ and $\alpha=0.05$. The maximum Mach number in this simulation is 0.246, remaining below the recommended limit to ensure acceptable compressibility errors [60].
@@ 3.3 Test 2: Left Ventricle with Thoracic Aorta
Figure 9: (a) Pressure and (b) flow rate on the inlet and planes A to D of the aorta domain. The results are obtained from the simulation using $k=4,\alpha=1.0$.
We examine the flow on different cross-sections of the aorta, namely the inlet and planes A to D (see Figure 10 for their locations). The plots in Figure 9 show that the pressure generally decreases as the flow moves downstream, while the flow rate attains a similar peak. The successive peaks for the inlet, planes A, B, C, and D, indicate the capture of wave propagation in the blood, which is observed in physiological humans [85]. Wave propagation is simulated because the LBM solves for a weakly compressible flow, allowing for slight changes in fluid density (see equation 13).
@@
Figure 10: (a) Flow velocity at the ejection peak ($t=0.88$ s) on (b) different cross-sections of the aorta. The heat map illustrates the magnitude of the normal velocity, while the arrows indicate the directions of the tangential velocity. The labels “O” and “I” indicate the outer and inner sides of the aorta, respectively.
@@
To gain more insights into the flow dynamics, we visualise the flow velocity on planes A to D. Figure 10 displays the magnitude of the normal velocity and the directions of the tangential velocity on these planes at the ejection peak ($t=0.88$ s). For all planes, the flow directions largely follow the shape of the vessels. For planes A and C, the flow accelerates as it flows over the bend, as suggested by the higher normal speed on the inner side. For plane B, the flow splits, with one direction leading upward towards the neck and the other downward into the descending aorta. These results are comparable to those observed at peak systole in studies examining aortic blood flow [84, 83].
@@
Figure 11: Displacement of heart muscles and blood pressure within the aorta during the ejection peak. The coupled model effectively captures the 3D structures of muscle contraction and vascular blood flow as well as their dynamical interactions.
@@
In addition, we visualise the displacement of heart muscles and blood pressure within the aorta during the ejection peak. As shown in Figure 11, the muscle contraction is more pronounced at the ventricle base and decreases towards the tip. Simultaneously, a high-pressure front originates from the inlet and propagates downstream. Around the outlets, the effects of the sponge layer are observed, where the pressure is lower than the surroundings. The coupled model effectively captures the 3D structures of muscle contraction and blood flow within the vessels as well as their dynamic interactions.
(a) Magnitude
(b) Difference from the baseline solution
Figure 12: (a) The magnitude of displacement of heart muscles at the end of the ejection phase ($t=0.97$ s) and (b) the difference compared with the baseline solution, i.e. solution from Test 2 minus that from Test 0).
@@
Furthermore, we compare the displacement of heart muscles at the end of the ejection phase with the baseline solution in Test 0. The results from these tests (Figures 6b and 12a) indicate a similar distribution of muscle displacement. However, the displacement in Test 2 ranges from 0.17 cm to 1.27 cm, whereas in Test 0 it ranges from 0.25 cm to 1.30 cm. As shown in Figure 12b, the differences between the two solutions generally decrease from the base to the tip and from the epicardium to the endocardium, with the greatest differences observed in the middle of the endocardium. These differences underscore the importance of accurately prescribing pressure values on the endocardium, which is determined by vascular blood flow.
@@
Since Alya and HemeLB are coupled in a staggered manner, the scalability of the coupled model is entirely determined by the scalability of the individual solvers, which has been extensively demonstrated in previous studies [41, 43]. The coupling scheme, which performs data exchange by reading and writing two small files (each containing only two numbers), runs sequentially. As a result, the time spent on coupling remains relatively constant and minimal. When Alya runs on two compute nodes and HemeLB on 64 nodes, the percentage of the time spent on coupling relative to the elapsed time per time step is about 4 % for Alya and 6 % for HemeLB (see A for more details).
@@ 5 Conclusion
We have successfully constructed a 3D model of the cardiovascular system by coupling the 3D electromechanical model of the heart in Alya with the 3D fluid mechanics model of vascular blood flow in HemeLB. Our file-based partitioned coupling scheme has been tested using idealised and realistic anatomies, demonstrating reliability and effectiveness. The implementation of a viscous sponge layer for blood flow and a low-pass filter at the coupling interface has proven effective in stabilising the coupled simulations. Notably, our coupled model predicts muscle displacement differently than the standalone heart model, highlighting the critical role of detailed vascular blood flow modelling in cardiac simulations. The coupling scheme is also efficient, requiring minimal additional computation time relative to advancing individual time steps in the heart and blood flow models.
Looking ahead, this coupled model can be further tested by applying it to a wide range of pathophysiological conditions, such as heart failure, hypertension, and congenital heart defects. Future work can also explore the integration of additional physiological systems, such as the respiratory and nervous systems, to enhance the fidelity of virtual human models and digital twins. Moreover, advancements in computational power and algorithms will likely allow for real-time simulations, thereby broadening the practical applications of these models in clinical decision-making and personalised medicine.
@@ Appendix A Computational Performance of the Coupling Scheme
We evaluate the computational performance of the coupling scheme by comparing the time spent on coupling with the time taken for one time step in each solver during the coupling phase.
We record the elapsed time of each time step in Alya and HemeLB and calculate their averages in the coupling phase. To determine the time spent on coupling, we measure the total time required to perform all procedures related to coupling, including the time for writing and reading data in files and the latency, and calculate its average. We perform these measurements using the problem outlined in Test 2.
@@
Figure 13: Average wall-clock time per time step and coupling time in Alya and HemeLB using various numbers of compute nodes. The results are shown for Alya using (a) one compute node and (b) two compute nodes with each compute node consisting of 128 CPU cores.
To understand the effects of parallelisation on the efficiency of the coupling scheme, we vary the number of CPU cores (or compute nodes) used by Alya and HemeLB. The results of the measurements are presented in Figure 13. While the average elapsed time for one time step decreases with the number of compute nodes used for HemeLB, the average time spent on coupling in Alya and HemeLB remains relatively constant between 0.01 s and 0.04 s. For the largest number of nodes tested, specifically two nodes for Alya and 64 nodes for HemeLB, the percentage of the time spent on coupling relative to the elapsed time per time step is about 4 % for Alya and 6 % for HemeLB.
@@
Sharp C. Y. Lo: Conceptualisation, Methodology, Software, Validation, Formal analysis, Investigation, Data curation, Writing — original draft, Visualisation. Alberto Zingaro: Conceptualisation, Methodology, Software, Validation, Formal analysis, Investigation, Writing — original draft, review & editing, Visualisation, Supervision. Jon W. S. McCullough: Conceptualisation, Methodology, Software, Funding acquisition. Xiao Xue: Software, Funding acquisition. Mariano Vázquez: Conceptualisation, Writing — review & editing, Supervision, Project administration, Funding acquisition. Peter V. Coveney: Conceptualisation, Writing — review & editing, Supervision, Project administration, Funding acquisition.

### v2

@@ 3 Results and Discussion
In this section, we present and analyse the results of the four simulation tests in Sections 3.13.4. Lastly, we assess the computational efficiency of the coupled model. All our simulations are performed on the ARCHER2 UK national supercomputer (https://www.archer2.ac.uk).
3.1 Test 0: Baseline Solutions for Left Ventricle and Thoracic Aorta
We run the standalone simulation of Alya for two heartbeats and present results from the second heartbeat, as the electromechanical model typically requires a few beats to reach a periodic solution. Figure 7a shows the time series of the left ventricular endocardial pressure and volume, with the ejection phase occurring from $t=0.842$ s to $t=0.973$ s. The endocardial pressure immediately before the ejection phase suggests a reference pressure $P_{\text{ref}}$ of 78 mmHg for the fluid domain in the subsequent tests.
(a) Left ventricular endocardial pressure and volume
(b) Aortic inflow rate
Figure 7: Results of Test 0. (a) Left ventricular endocardial pressure and volume from the standalone Alya simulation. The endocardial pressure immediately before the ejection phase suggests a reference pressure $P_{\text{ref}}$ of 78 mmHg for the fluid domain in the subsequent tests. (b) Inlet flow rate profile prescribed in the standalone HemeLB simulation, obtained as the negative derivative of the ventricular volume time series during the ejection phase with a prepended warm-up period (see equation 25d).
The ventricular volume in the ejection phase is used to derive the inlet flow rate profile for the fluid domain, as shown in Figure 7b, following equation 25d. To ensure a smooth transition, a warm-up period of 0.025 s is prepended to the inflow profile, during which the flow rate increases steadily to its initial value. We run the standalone simulation of HemeLB with this inlet profile. Figures 14a and 15a show the magnitude of heart muscle displacement and the time-averaged wall shear stress in the aorta, respectively, for comparison with Test 2.
3.2 Test 1: Coupled Simulations of Left Ventricle and Cylinder
Here we present the results of the coupled simulations using the left ventricle and cylinder geometries. First of all, we assess the reliability of the coupling scheme. We perform the coupled simulations using different time step sizes and coupling frequencies in HemeLB while keeping them fixed in Alya. These simulations are completed at the specified simulation times, showing that our implementation of the coupling algorithm is reliable.
Figure 8: Pressure and flow rate values at the coupling interface from Alya and HemeLB, demonstrating consistency between the two codes during coupling. The results are obtained from the coupled simulation with $k=4$ in Test 1.
Second, we check the consistency of the exchanged data. During the coupling phase, the pressure and flow rate values at the coupling interface from Alya and HemeLB show excellent agreement. As an example, the outputs for $k=4$ are presented in Figure 8. This suggests that the exchanged data are consistent in the two solvers and that our coupling algorithm is correctly implemented.
@@ 3.2 Test 1: Coupled Simulations of Left Ventricle and Cylinder
Figure 9: (a) Pressure and (b) flow rate on the inlet, middle, and outlet planes of the cylinder domain in Test 1. These variables remain bounded and free of spurious oscillations, demonstrating the stability of the coupled framework.
Third, we evaluate the stability of the coupled simulations. Instabilities are more likely to arise in the fluid dynamics simulations because the flow variables are highly sensitive to numerical oscillations. Therefore, we focus on assessing stability through the pressure and flow rate in the HemeLB simulations. We find that these variables remain bounded and free of spurious oscillations at the inlet, middle, and outlet of the fluid domain throughout the simulations. As an example, Figure 9 presents the results for $k=4$. The same is observed for the electromechanical solutions of Alya. These results confirm the stability of the coupled framework.
@@
Figure 10: Convergence of errors in (a) pressure ($\varepsilon_{P}$) and (b) flow rate ($\varepsilon_{Q}$) in the coupled Alya-HemeLB simulation of Test 1. The pressure and flow rate are measured at the inlet at the ejection peak (0.865 s) across four resolutions. Their absolute errors are computed relative to the solutions on the finest grid ($\Delta x^{\mathcal{H}}=9.375\cdot 10^{-5}$ m and $\Delta t^{\mathcal{H}}=1.40625\cdot 10^{-5}$ s). The observed convergence rates are approximately one for the pressure and two for the flow rate.
Fourth, we assess the impact of grid resolution on the solutions of the coupled model. The absolute errors in pressure and flow rate are shown in Figure 10. The observed convergence rates are approximately one for the pressure and two for the flow rate. We note that the differences in discretisation schemes between the lattice Boltzmann and finite element methods may affect error propagation, potentially influencing the observed convergence rates. Despite this, the second-order convergence of the flow rate suggests that the overall flow dynamics remain well captured across different grid resolutions.
3.3 Test 2: A Coupled Simulation of Left Ventricle and Thoracic Aorta
Here we present the results of the coupled simulations using the left ventricle and thoracic aorta geometries. Initially, the coupled simulations become unstable at the tested coupling frequencies ($k=8,10,12,14,16,20$). However, applying the low-pass filter with an appropriate smoothing factor ($\alpha$), all these simulations become stable. This demonstrates the effectiveness of the low-pass filter in stabilising the coupled model for this specific system. In the following, we look into the simulation using $k=16$ and $\alpha=0.05$. The maximum Mach number in this simulation is 0.148, remaining below the recommended limit to ensure acceptable compressibility errors [60].
(a) Pressure
(b) Flow rate
@@ 3.3 Test 2: A Coupled Simulation of Left Ventricle and Thoracic Aorta
Figure 11: (a) Pressure and (b) flow rate at the inlet and planes A to D (see Figure 12 for their locations) of the aorta in Test 2.
To characterise the flow behaviour, we analyse the pressure and flow rate at the aortic inlet and cross-sectional planes A to D (see Figure 12 for their locations). Figure 11 presents the acquired time series. Both variables show amplitudes and waveforms consistent with in vivo measurements [92, 93, 94, 95]. The successive peaks from the inlet to plane D reflect wave propagation effects similar to those observed in human physiology [96]. Wave propagation is simulated because the LBM solves for a weakly compressible flow, allowing for slight changes in fluid density (see equation 15).
@@
Figure 12: (a) Flow velocity at the ejection peak ($t=0.88$ s) on (b) different cross-sections of the aorta in Test 2. The heat map illustrates the magnitude of the normal velocity, while the arrows indicate the directions of the tangential velocity. The labels “O” and “I” indicate the outer and inner sides of the aorta, respectively.
@@
For a more detailed understanding of the flow dynamics, we visualise the velocity field on planes A to D. Figure 12 shows the magnitude of the normal velocity and the directions of the tangential velocity at the ejection peak ($t=0.88$ s). Across all planes, the flow directions align with the vessel morphology. On planes A and C, the flow accelerates along the inner bend, evidenced by the higher normal velocities. Plane B exhibits flow splitting, with one component directed upward towards the neck vessels and the other downward into the descending aorta. These results are comparable to those observed at the peak systole in studies examining aortic blood flow [97, 98].
@@
Figure 13: Displacement of heart muscles and blood pressure within the aorta during the coupling phase in Test 2. The coupled model effectively captures the 3D heart contraction and vascular blood flow as well as their dynamical interactions.
@@
To explore how the heart muscles and vascular blood flow interact, we visualise the displacement of heart muscles and blood pressure within the aorta during the coupling phase. As shown in Figure 13, the muscle contraction is more pronounced at the ventricle base and decreases towards the apex. Simultaneously, a high-pressure front originates from the inlet and propagates downstream. Around the outlets, the effects of the sponge layer are observed, where the pressure is lower than the surroundings. The coupled model effectively captures the 3D heart contraction and vascular blood flow within the vessels as well as their dynamical interactions.
Figure 14: Magnitude of heart muscle displacement at the end of the ejection phase ($t=0.97$ s) for (a) the standalone Alya simulation in Test 0 and (b) the coupled Alya-HemeLB simulation in Test 2, and (c) the percentage difference of Test 2 relative to Test 0.
@@
To assess the impact of coupling on cardiac mechanics, we compare the displacement of the heart muscles at the end of the ejection phase between Test 0 and Test 2. Figure 14a shows the displacement from the standalone Alya simulation (Test 0), while Figure 14b presents the results from the coupled simulation (Test 2). Both exhibit a similar distribution of muscle displacement; however, there are notable differences in magnitude. As shown in Figure 14c, the percentage difference between the two solutions increases from the base to the apex and from the epicardium to the endocardium, peaking at 22 % close to the apex at the endocardium. These differences underscore the importance of accurately prescribing endocardial pressure, which is determined by vascular blood flow.
Figure 15: Time-averaged wall shear stress (TAWSS) in the aorta during the ejection phase for (a) the standalone HemeLB simulation in Test 0, (b) the coupled Alya-HemeLB simulation in Test 2, and (c) the percentage difference of Test 2 relative to Test 0.
Similarly, we evaluate the influence of coupling on the TAWSS in the aorta during the ejection phase between Test 0 and Test 2, as shown in Figure 15. While the overall distribution of TAWSS in the standalone HemeLB simulation (Test 0) is similar to that in the coupled simulation (Test 2), notable differences in magnitude are observed. As shown in Figure 15c, the percentage difference is more pronounced in the upstream region, reaching up to 20 % in parts of the ascending aorta. These differences highlight the critical role of accurately prescribing the inlet profile, which is governed by the cardiac output resulting from the coupling with the ventricular electromechanics.
@@
Since Alya and HemeLB are coupled in a staggered manner, the scalability of the coupled model is entirely determined by the scalability of the individual solvers, which has been extensively demonstrated in previous studies [41, 43]. The coupling scheme, which performs data exchange by reading and writing two small files (each containing only two numbers), runs sequentially. As a result, the time spent on coupling remains relatively constant and minimal. When Alya runs on two compute nodes and HemeLB on 64 nodes, the percentage of the time spent on coupling relative to the elapsed time per coupling iteration is about 4 % for Alya and 6 % for HemeLB (see A for more details).
3.4 Test 3: Impact of Myocardial Scarring on Downstream Vascular Flow
Here we present the results of the coupled simulations with cardiac scarring and analyse the impact of the scarring on downstream vascular flow.
Figure 16: Left ventricular contraction at the ejection peak in Test 3, with myocardial scars of varying sizes (bounded by grey lines): 0.5 cm (left, peak at 0.88 s), 1.0 cm (middle, 0.88 s), and 2.5 cm (right, 0.93 s). The transparent overlay shows the left ventricular anatomy in its initial configuration.
Figure 16 shows the contraction of the left ventricle with myocardial scars of different sizes at the ejection peak. As the scar size increases, noticeable changes in the contraction pattern emerge. To better quantify these differences, Figure 17 presents the time series of the pressure and flow rate at the coupling interface for different scar sizes, alongside the scar-free case from Test 2 for comparison. As the scar size increases, both pressure and flow decrease, and the onset of the ejection phase is delayed. These deviations from the scar-free case become more pronounced with larger scars, highlighting the influence of myocardial damage on cardiac function and haemodynamics.
(a) Pressure
(b) Flow rate
Figure 17: (a) Pressure and (b) flow rate at the coupling interface in the Alya-HemeLB simulations in Test 3, each using the same heart model but with scars of varying sizes, coupled with the thoracic aorta model. Compared to the scar-free case (Test 2), both the pressure and the flow rate decrease more significantly as the scar size increases, and the ejection phase starts later.
To further investigate the haemodynamic impact of myocardial scarring, we examine the distribution of TAWSS in the aorta during the coupling phase. Figure 18 shows the TAWSS distributions for different scar sizes, along with the percentage difference relative to the scar-free case from Test 2 (Figure 15b). As the scar size increases, deviations in the TAWSS distribution from the scar-free case become more noticeable. The maximum difference in the aorta increases from 3.0 % in the 0.5 cm scar case to 25 % in the 1.0 cm scar case and reaches up to 87 % in the 2.5 cm scar case. These results highlight the substantial impact of myocardial scarring on aortic haemodynamics, which may have important implications for vascular disease development, as abnormal TAWSS is linked to conditions such as aneurysms, atherosclerosis, and thrombosis [88, 89].
Figure 18: Time-averaged wall shear stress (TAWSS) in the aorta during the ejection phase of the coupled simulations in Test 3, using heart models with myocardial scars of different sizes: (a) 0.5 cm, (b) 1.0 cm, (c) 2.5 cm. The percentage difference of (a), (b), and (c) relative to Test 2 in Figure 15b is shown in (d), (e), and (f).
@@ 4 Limitations
A fourth limitation is the assumption of a rigid aorta, which neglects its natural compliance. In reality, the aortic cross-section undergoes significant changes throughout the cardiac cycle, as demonstrated in both clinical [103, 104, 105] and computational studies, where fluid-structure interaction models are developed [92, 106, 107, 108, 109]. This limitation affects the computed flow velocity and pressure, as arterial compliance plays a crucial role in dampening pressure waves and modulating blood flow dynamics. Within a lattice Boltzmann framework, fluid–structure interactions can be modelled by coupling the fluid dynamics with a finite element representation of the aortic wall mechanics, for instance, through the immersed boundary method [32, 110].
@@ 5 Conclusion
We have successfully constructed a 3D model of the cardiovascular system by coupling the 3D electromechanical model of the heart in Alya with the 3D fluid mechanics model of vascular blood flow in HemeLB. Our file-based partitioned coupling scheme has been tested using idealised and realistic anatomies, demonstrating its reliability and effectiveness. The implementation of a viscous sponge layer for blood flow and a low-pass filter at the coupling interface has proven effective in stabilising the coupled simulations. The coupling scheme is also efficient, requiring minimal additional computation time relative to advancing individual time steps in the heart and blood flow models.
Notably, our coupled model predicts muscle displacement and aortic wall shear stress differently than the standalone heart and blood flow models, highlighting the critical importance of coupling between cardiac and vascular dynamics to capture their intricate interactions in cardiovascular simulations. Muscle displacement shows peak differences of up to 22 % at the endocardium, while TAWSS in the ascending aorta differs by as much as 20 %. Furthermore, our simulations highlight the model’s ability to capture key features of cardiac disease by assessing the impact of myocardial scarring on downstream vascular flow. This is made possible by the coupled interaction between the cardiac and vascular dynamics. As scar severity increases, both cardiac function and vascular haemodynamics deteriorate, with variations in TAWSS reaching 87 % in the most severe case.
Looking ahead, this coupled model can be further tested by applying it to a wide range of pathophysiological conditions, such as heart failure, hypertension, and congenital heart defects. Future work can also explore the integration of additional physiological systems, such as the respiratory and nervous systems, to enhance the fidelity of virtual human models and digital twins. Moreover, advancements in computational power and algorithms will likely allow for real-time simulations, thereby broadening the practical applications of these models in clinical decision making and personalised medicine.
@@ Appendix A Computational Performance of the Coupling Scheme
We evaluate the computational performance of the coupling scheme by comparing the time spent on coupling with the time taken for one coupling step in each solver during the coupling phase.
We record the elapsed time of each coupling iteration in Alya and HemeLB and calculate their averages in the coupling phase. Here, one coupling iteration is equivalent to one time step in Alya and 16 time steps in HemeLB (since $k=16$ in Test 2). To determine the time spent on coupling, we measure the total time required to perform all procedures related to coupling, including the time for writing and reading data in files and the latency, and calculate its average. We perform these measurements using the problem outlined in Test 2.
@@
Figure 19: Average elapsed time and coupling time in Alya and HemeLB per coupling iteration using various numbers of compute nodes. The results are shown for Alya using (a) one compute node and (b) two compute nodes with each compute node consisting of 128 CPU cores.
To understand the effects of parallelisation on the efficiency of the coupling scheme, we vary the number of CPU cores (or compute nodes) used by Alya and HemeLB. The results of the measurements are presented in Figure 19. While the average elapsed time for one coupling iteration decreases with the number of compute nodes used for HemeLB, the average time spent on coupling in Alya and HemeLB remains relatively constant between 0.01 s and 0.04 s. For the largest number of nodes tested, specifically two nodes for Alya and 64 nodes for HemeLB, the percentage of the time spent on coupling relative to the elapsed time per coupling iteration is about 4 % for Alya and 6 % for HemeLB.
In our current setup, the computational overhead from coupling is minimal because only a small amount of data is exchanged at each time step. Under these conditions, the file-based partitioned coupling approach remains efficient. However, in cases where large amounts of data need to be exchanged, such as full 3D fields in fluid–structure interaction or electromechanics simulations, using a dedicated coupling framework like MUSCLE3 [112] or a monolithic software approach [5, 37, 35, 7, 8] may be more efficient. These approaches avoid repeated file I/O and allow for better exploitation of parallelism within the solver. Determining the point at which file-based coupling becomes inefficient depends on multiple factors, including data volume, exchange frequency, and solver architecture. Beyond performance, long-term maintainability and the planned longevity of the coupling framework also play a crucial role in selecting the most suitable approach. These computational and practical considerations together shape the choice of coupling strategy and represent an interesting direction for further investigation.
@@
Sharp C. Y. Lo: Conceptualisation, Methodology, Software, Validation, Formal analysis, Investigation, Data curation, Writing — original draft, Visualisation. Alberto Zingaro: Conceptualisation, Methodology, Software, Validation, Formal analysis, Investigation, Writing — original draft, review & editing, Visualisation, Supervision. Jon W. S. McCullough: Conceptualisation, Methodology, Software, Writing — review & editing, Funding acquisition. Xiao Xue: Software, Funding acquisition. Pablo Gonzalez-Martin: Methodology, Software. Balint Joo: Software, Writing — review & editing. Mariano Vázquez: Conceptualisation, Writing — review & editing, Supervision, Project administration, Funding acquisition. Peter V. Coveney: Conceptualisation, Writing — review & editing, Supervision, Project administration, Funding acquisition.